# Lepton Flavor-Violating Higgs Decays Mediated by Ultralight Gauge Boson

### v1

@@ Abstract
We present an analysis of the lepton-flavor violating decay of the Higgs boson mediated by an ultralight gauge boson, $\chi$. Our analysis matches a model generating the lepton flavor-violating interaction ¯$\bar{\ell}_{i}\ell_{j}\chi$ at tree level with an effective field theory, safeguarding the $\chi$-boson mass as it tends to zero. By utilizing the upper bounds on $H\to\ell_{i}\bar{\ell}_{j}$ decays from CMS and ATLAS Collaborations, we establish an indirect upper limit on the nonstandard decay $H\to\ell_{i}\bar{\ell}_{j}\chi$. The analysis encompasses various observables such as the lepton energy spectrum, Dalitz plot distribution, and Lepton Charge and Forward-Backward Asymmetries.
@@ 1 Introduction
Lepton-flavor violating (LFV) decays $H\rightarrow\mu e$ [^1] In this manuscript, the notation $H\to\mu e$ refers to $H\to\mu\bar{e}$ or $H\to\bar{\mu}e$. The same convention applies to the other cLFV modes., $H\rightarrow\tau e$, or $H\rightarrow\tau\mu$ are strictly forbidden within the confines of the SM framework and effectively suppressed in its minimal extension [14]. However, they find theoretical justification in BSM theories, where LFV interaction can arise through off-diagonal LFV Yukawa couplings, which facilitate the interaction of the Higgs boson with leptons of differing flavors. These processes are primarily induced via virtual boson exchange mechanisms [15, 16, 17, 18, 19], often occurring within extended Higgs sectors, or involving several scalar doublets [20, 21, 22, 23, 24, 25, 26]. LHC Collaborations have undertaken experimental efforts to probe Higgs decays [27]. Specifically, the ATLAS and CMS Collaborations have conducted extensive measurements for the SM decay modes $H\rightarrow\tau\bar{\tau}$ and $H\rightarrow\mu\bar{\mu}$, alongside establishing upper bounds for $H\rightarrow e\bar{e}$ and all potentially cLFV channels of $H\rightarrow\ell_{i}\bar{\ell}_{j}$ decay. The outcomes of these searches are detailed in Table 1.
@@ 2 Effective Lagrangian description
In the letter [30], one of the authors proposes two models enabling LFV interactions via a gauge boson, denoted as $\chi$. Their focus lies on an ultralight gauge boson, which ensures finiteness in the massless $\chi$-boson limit. Although the initial paper confines the models to two generations, expanding them to encompass three generations is straightforward [31]. We are especially intrigued by the tree-level model that induces cLFV transitions at the tree level.
@@
Table 2: Spins and charges under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{\chi}$ model that describes LFV transitions at tree level. All fields are assumed to be singlets under $SU(3)_{C}$.
@@
where it can be deduced that the mixing angles $\theta_{L}\,,~{}\theta_{R}$ and the masses of the particle can be written as a function of $y_{jk}$, $v_{jk}$ and $q_{\phi_{jk}}$, see Appendix A.
@@
By simplicity, if we assumed that the Yukawa couplings satisfy $y_{22}>>y_{11}>>y_{12},y_{21}$ and all $v_{jk}=v$, so $q_{\phi_{jk}}=Q$. The relevant parameters of the tree-level model after the spontaneous breaking of the gauge symmetries are:
@@ 3 LFV Higgs decays
In the Standard Model, the flavor-conserving interaction ¯$\bar{\ell}_{i}\ell_{i}H$ emerges, while our effective Lagrangian, as outlined in Eq. (9), describes transitions for all lepton interactions, encompassing scenarios with flavor-conserving interaction $\ell_{i}=\ell_{j}$ and flavor-violating interactions $\ell_{i}\neq\ell_{j}$. Using this effective Lagrangian, we can induce LFV processes $H\to\ell_{i}\bar{\ell}_{j}$ at the loop level mediated by $\chi$, as well as $H\to\ell_{i}\bar{\ell}_{j}\chi$ at the tree level. The forthcoming two subsections will present the analytical and phenomenological outcomes of these decay processes.
@@ 3.1 Higgs LFV decays with $\boldsymbol{\chi}$ off-shell
$\simeq\frac{m_{\chi}^{2}}{m_{\ell_{i}^{k}}^{2}m_{\ell_{j}^{k}}^{2}}$ $\frac{M_{H}m_{\ell_{k}}^{2}}{4\pi~{}\Gamma_{H}~{}v^{2}}\bigl{[}\lvert c_{jk}^{v}c_{ik}^{a}-c_{ik}^{v}c_{jk}^{a}\rvert^{2}+\lvert c_{jk}^{v}c_{ik}^{v}-c_{ik}^{a}c_{jk}^{a}\rvert^{2}\bigr{]}~{}\bigl{\lvert}\mathcal{F}(m_{\ell_{i}}\,,m_{\ell_{j}}\,,m_{\ell_{k}})\bigr{\rvert}^{2}\,,$
@@
$\mathcal{F}(m_{\ell_{i}}\,,m_{\ell_{j}}\,,m_{\ell_{k}})=$ $\frac{1}{m_{\chi}^{2}}{\rm A0}\left[m_{\ell_{k}}^{2}\right]-\left(\frac{M_{H}^{2}}{2m_{\chi}^{2}}-3\right){\rm B0}\left[M_{H}^{2}\,,m_{\ell_{k}}^{2}\,,m_{\ell_{k}}^{2}\right]-3\Bigl{(}{\rm B0}\left[m_{\ell_{i}}^{2}\,,m_{\ell_{k}}^{2}\,,m_{\chi}^{2}\right]+$
[…]
A0, B0, and C0 denote the Passarino-Veltman integrals, and $\Lambda$, encoded within these integrals, represents the cutoff energy scale associated with our EFT framework. The function $\mathcal{F}(m_{\ell_{i}},m_{\ell_{j}},m_{\ell_{k}})$ depends on the masses involved in the process, including the Higgs mass, the lepton masses, and the $\chi$-boson mass, as well as $\Lambda$. It is important to note that the $\chi$-boson mass in the denominator of $\mathcal{F}(m_{\ell_{i}},m_{\ell_{j}},m_{\ell_{k}})$ is canceled by the $m_{\chi}^{2}$ term in the expression for ${\rm BR}(H\to\ell_{i}\ell_{j})$ in Eq. (11), thus ensuring finitude in the massless limit of $m_{\chi}$.
From Eq. (11), it is evident that the branching ratio ${\rm BR}(H\to\ell_{i}\ell_{j})$ relies on six free parameters: $\Lambda$, $m_{\chi}$, and the coefficients $c_{ik}^{v}$, $c_{jk}^{v}$, $c_{ik}^{a}$, $c_{jk}^{a}$. Please be aware that when $c_{ik}^{v}$ and $c_{ik}^{a}$ are equal, the model constraints dictate that $\Gamma(H\to\ell_{i}\bar{\ell}_{j})$ vanishes identically. This arises from the fact that if $c_{ik}^{v}=c_{ik}^{a}$, the charges under $U(1)_{\chi}$ are generation-independent, precluding any tree-level LFV interaction. Our focus lies in a scenario where $\chi$ is a light boson, and the decay $\mu\to e\chi$ is viable, implying $m_{\chi}\leq m_{\mu}$. Additionally, our attention is directed toward low-energy processes, thus $\Lambda$ remains below the electroweak scale.
@@
In Figures 2, we present the branching ratios ${\rm BR}(H\to\ell_{i}\ell_{j})$ as a function of $m_{\chi}$ for the three cLFV channels: $H\to\tau\mu$, $H\to\tau e$, and $H\to\mu e$. For our numerical analysis, we fix $\Lambda=M_{H}$, consider $m_{\chi}$ ranging from 00 to $m_{\mu}$, and utilize the experimental total Higgs decay width, $\Gamma_{H}=3.7^{~{}+1.9}_{~{}-1.4}$ MeV [32]. To ensure clarity and ease of interpretation, we adopt the simplifying assumptions $c_{jk}^{a}=0$, thereby suppressing the axial contribution. Suitable values are assigned to $c_{jk}^{v}$, as indicated in each figure. The red line corresponds to the scenario where all three lepton contributions are active within the loop, i.e., $\tau$, $\mu$, and $e$. Conversely, when only a single lepton contribution is activated into the loop, we represent it with purple, green, and blue lines for $e$, $\mu$, and $\tau$, respectively. To contextualize our findings, the grey line denotes the current upper limit, as tabulated in Table 1. Dashed lines are employed to address cases of line overlap, ensuring clarity in the visualization of results.
@@
Our observations demonstrate distinct dominant contributions for each decay process, as depicted in Figures 2(a), 2(b), and 2(c). For the decay $H\to\tau\mu$, the diagrams featuring $\tau$ and $\mu$ within the loop yield the dominant contributions. In the case of $H\to\tau e$, all three diagrams, with $\tau$, $\mu$, and $e$ in the loop, exhibit comparable significance. Conversely, for $H\to\mu e$, the dominant diagrams involve $\mu$ and $e$. This behavior is explained by Eq. (11), which states that $\mathrm{BR}(H\to\ell_{i}\ell_{j})\propto\frac{m_{\ell_{k}}}{m_{\ell_{i}}m_{\ell_{j}}}$. The masses in the denominator result from the relationship between the effective couplings $f_{jk}$ and $g_{jk}$ with the coefficients $c_{jk}^{v}$ and $c_{jk}^{a}$ as described in Eq. (10).
@@
By employing the upper limits provided for ${\rm BR}(H\to e\mu\,,e\tau\,,\mu\tau)$ and the flavor-conserving processes ${\rm BR}(H\to ee\,,\mu\mu\,,\tau\tau)$, which are detailed in Table 1, we can discern constraints on the individual coefficients $|c_{ik}^{v(a)}|$. Under the assumption that all $|c_{ik}^{v}|\neq 0~{}{\rm and}~{}|c_{ik}^{a}|=0$, and with the parameters $m_{\chi}=m_{\mu}/2$ and $\Lambda=M_{H}$ fixed, we derive restrictions for all $|c_{ik}^{v}|$.
@@
It is important to emphasize that these constraint regions remain valid across different values of $m_{\chi}$, as we have shown that ${\rm BR}(H\to\ell_{i}\ell_{j})$ exhibits minimal dependence on $m_{\chi}$. Furthermore, under the assumption that $c_{ik}^{a}\neq 0$ and $c_{ik}^{v}=0$, with consistent values for $m_{\chi}$ and $\Lambda$, the resultant constraint regions would remain consistent. This limit on the coefficients $c_{ii}^{v}$ can, in turn, be translated into constraints on the Yukawa couplings, $U(1)_{\chi}$ charges and gauge coupling, and expectation values of the scalar doublets $\phi_{ij}$ in the tree-level model, ensuring the correct couplings with the Higgs are reproduced. Let us finish this section, noting that based on the analysis performed in [30] for the tree-level model, it is anticipated that the lepton three-body decays $\mu\to 3e$ and $\tau\to 3\ell$ (in six possible combinations when $\ell=e$ or $\mu$) would provide the most stringent constraints on the coefficients $c_{jk}^{v(a)}$. However, the primary focus of this work is not on deriving the optimal constraints for the coefficients $c_{jk}^{v(a)}$, but rather on examining the lepton-flavor violating decay mediated by an ultralight boson within the context of Higgs decays.

### v2

@@ Abstract
We present an analysis of the lepton-flavor violating decay of the Higgs boson mediated by an ultralight gauge boson, $\chi$. Our analysis matches a model generating the lepton flavor-violating interaction ¯$\bar{\ell}_{i}\ell_{j}\chi$ at tree level with an effective field theory, safeguarding a physical massless $\chi$-boson limit of the observables. By utilizing the upper bounds on $H\to\ell_{i}\bar{\ell}_{j}$ decays from CMS and ATLAS Collaborations, we establish an indirect upper limit on the nonstandard decay $H\to\ell_{i}\bar{\ell}_{j}\chi$. The analysis encompasses various observables such as the lepton energy spectrum, Dalitz plot distribution, and Lepton Charge and Forward-Backward Asymmetries.
@@ 1 Introduction
Lepton-flavor violating (LFV) decays $H\rightarrow\mu e$ [^1] In this manuscript, the notation $H\to\mu e$ refers to $H\to\mu\bar{e}$ and $H\to\bar{\mu}e$. The same convention applies to the other cLFV modes., $H\rightarrow\tau e$, or $H\rightarrow\tau\mu$ are strictly forbidden within the confines of the SM framework and effectively suppressed in its minimal extension [14]. However, they find theoretical justification in BSM theories, where LFV interaction can arise through off-diagonal LFV Yukawa couplings, which facilitate the interaction of the Higgs boson with leptons of differing flavors. These processes are primarily induced via virtual boson exchange mechanisms [15, 16, 17, 18, 19], often occurring within extended Higgs sectors, or involving several scalar doublets [20, 21, 22, 23, 24, 25, 26]. LHC Collaborations have undertaken experimental efforts to probe Higgs decays [27, 28, 29]. Specifically, the ATLAS and CMS Collaborations have conducted extensive measurements for the SM decay modes $H\rightarrow\tau\bar{\tau}$ and $H\rightarrow\mu\bar{\mu}$, alongside establishing upper bounds for $H\rightarrow e\bar{e}$ and all potentially cLFV channels of $H\rightarrow\ell_{i}\bar{\ell}_{j}$ decays. The outcomes of these searches are detailed in Table 1.
@@ 2 Effective Lagrangian description
In the letter [30], one of the authors proposes two models enabling LFV interactions via a gauge boson, denoted as $\chi$. Their focus lies on an ultralight gauge boson, which ensures the finiteness of the observables in the massless $\chi$-boson limit. Although the initial paper confines the models to two generations, expanding them to encompass three generations is straightforward [31]. We are especially intrigued by the tree-level model that induces cLFV transitions at the tree level.
@@
Table 2: Spins and charges under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{\chi}$ model describing LFV transitions at tree level in the two generation case. All fields are assumed to be singlets under $SU(3)_{C}$.
@@
where it can be deduced that the mixing angles $\theta_{L}\,,~{}\theta_{R}$ and the masses of the particles can be written as a function of $y_{jk}$, $v_{jk}$ and $q_{\phi_{jk}}$, see Appendix A. Evidently, when the $U(1)_{\chi}$ charges are generation-independent, lepton flavor is exactly conserved within the model. Likewise, if the interaction eigenstates coincide with the mass eigenstates, tree-level flavor-changing interactions are absent (see Eq. (22)). This alignment of eigenstates can occur under specific conditions, such as particular configurations of the vacuum expectation values of the fields $\phi_{ij}$, for example, when $v_{ii}=0$ but $v_{ij}\neq 0$. This mechanism holds not only for two generations but can also be extended to three generations [31].
@@
For simplicity, we assume for the moment that the Yukawa couplings satisfy $y_{22}>>y_{11}>>y_{12},y_{21}$ and all $v_{jk}=v$, so $q_{\phi_{jk}}=Q$. The relevant parameters of the tree-level model after the spontaneous breaking of the gauge symmetries are:
@@
In this work, we focus on the scenario where $\chi$ is (ultra)light, specifically $m_{\chi}<m_{\mu}$ (with $m_{\chi}\ll m_{\mu}$ in the ultralight case). The effective parameters can be constrained to ensure the correct muon mass $m_{\mu}\sim 105$ MeV by setting appropriate limits on the Yukawa couplings, $U(1)_{\chi}$-charges, and the gauge coupling $g_{\chi}$, along with the vacuum expectation values of the scalar doublets $\phi_{ij}$. To achieve a light $\chi$-boson, one may select a small $g_{\chi}$, small $q_{\phi_{ij}}$, or both, leading to an ultralight $\chi$-boson while preserving the observed lepton mass spectrum.
@@ 3 LFV Higgs decays
In the Standard Model, the flavor-conserving interaction ¯$\bar{\ell}_{i}\ell_{i}H$ emerges, while our effective Lagrangian, as outlined in Eq. (9), describes transitions for all lepton interactions, encompassing scenarios with flavor-conserving interactions $\ell_{i}=\ell_{j}$ and flavor-violating interactions $\ell_{i}\neq\ell_{j}$. Using this effective Lagrangian, we can induce LFV processes $H\to\ell_{i}\bar{\ell}_{j}$ at the loop level mediated by $\chi$, as well as $H\to\ell_{i}\bar{\ell}_{j}\chi$ at the tree level. The forthcoming two subsections will present the analytical and phenomenological outcomes of these decay processes.
@@ 3.1 Higgs LFV decays with $\boldsymbol{\chi}$ off-shell
$\simeq\frac{m_{\chi}^{4}}{m_{\ell_{i}^{k}}^{2}m_{\ell_{j}^{k}}^{2}}$ $\frac{M_{H}m_{\ell_{k}}^{2}}{4\pi~{}\Gamma_{H}~{}v^{2}}\bigl{[}\lvert c_{jk}^{v}c_{ik}^{a}-c_{ik}^{v}c_{jk}^{a}\rvert^{2}+\lvert c_{jk}^{v}c_{ik}^{v}-c_{ik}^{a}c_{jk}^{a}\rvert^{2}\bigr{]}~{}\bigl{\lvert}\mathcal{F}_{{\rm ren}}(m_{\ell_{i}}\,,m_{\ell_{j}}\,,m_{\ell_{k}})\bigr{\rvert}^{2}\,,$
@@
$\mathcal{F}(m_{\ell_{i}}\,,m_{\ell_{j}}\,,m_{\ell_{k}})=$ $\frac{{\rm A_{0}}\left[m_{\ell_{k}}^{2}\right]}{m_{\chi}^{2}}-\left(\frac{M_{H}^{2}}{2m_{\chi}^{2}}-\frac{2m_{\ell_{k}}^{2}}{m_{\chi}^{2}}-3\right){\rm B_{0}}\left[M_{H}^{2}\,,m_{\ell_{k}}^{2}\,,m_{\ell_{k}}^{2}\right]-3\Bigl{(}{\rm B_{0}}\left[m_{\ell_{i}}^{2}\,,m_{\ell_{k}}^{2}\,,m_{\chi}^{2}\right]$
$+{\rm B_{0}}$ $\left[m_{\ell_{j}}^{2}\,,m_{\ell_{k}}^{2}\,,m_{\chi}^{2}\right]-\left(M_{H}^{2}+m_{\chi}^{2}\right){\rm C_{0}}\left[M_{H}^{2}\,,m_{\ell_{i}}^{2}\,,m_{\ell_{j}}^{2}\,,m_{\ell_{k}}^{2}\,,m_{\ell_{k}}^{2}\,,m_{\chi}^{2}\right]\Bigr{)}+2\,.$ (12)
A0, B0, and C0 denote the Passarino-Veltman integrals. The function $\mathcal{F}(m_{\ell_{i}},m_{\ell_{j}},m_{\ell_{k}})$ depends on the masses involved in the process, including the Higgs mass, the lepton masses, and the $\chi$-boson mass. The loop function in Eq. (3.1) exhibits UV divergences, and it can be expressed as the sum of a divergent and a finite (renormalized) part, denoted as $\mathcal{F}=\mathcal{F}_{\text{ren}}+\mathcal{F}_{\text{div}}$ where $\mathcal{F}_{\text{ren}}$ refers to the renormalized contribution, and $\mathcal{F}_{\text{div}}$ captures the divergent component (see Appendix B for details). It is important to note that the $\chi$-boson mass in the denominator of $\mathcal{F}(m_{\ell_{i}},m_{\ell_{j}},m_{\ell_{k}})$ is canceled by the $m_{\chi}^{2}$ term in the expression for ${\rm BR}(H\to\ell_{i}\ell_{j})$ in Eq. (11), thus ensuring finitude in the massless limit of $m_{\chi}$.
From Eq. (11), it is evident that the branching ratio ${\rm BR}(H\to\ell_{i}\ell_{j})$ relies on five free parameters: $m_{\chi}$ and the coefficients $c_{ik}^{v}$, $c_{jk}^{v}$, $c_{ik}^{a}$, $c_{jk}^{a}$. Please be aware that when $c_{ik}^{v}$ and $c_{ik}^{a}$ are equal, the model constraints dictate that $\Gamma(H\to\ell_{i}\bar{\ell}_{j})$ vanishes identically. This arises from the fact that if $c_{ik}^{v}=c_{ik}^{a}$, the charges under $U(1)_{\chi}$ are generation-independent, precluding any tree-level LFV interaction. Our focus lies in a scenario where $\chi$ is a light boson, and the decay $\mu\to e\chi$ is viable, implying $m_{\chi}\leq m_{\mu}$. Additionally, our attention is directed toward low-energy processes, thus $\Lambda$ remains below the electroweak scale.
@@
In Figures 2, we present the branching ratios ${\rm BR}(H\to\ell_{i}\ell_{j})$ as a function of $m_{\chi}$ for the three cLFV channels: $H\to\tau\mu$, $H\to\tau e$, and $H\to\mu e$. For our numerical analysis, we consider $m_{\chi}$ ranging from 00 to $m_{\mu}$, and utilize the experimental total Higgs decay width, $\Gamma_{H}=\left(3.7^{~{}+1.9}_{~{}-1.4}\right)$ MeV [32]. To ensure clarity and ease of interpretation, we adopt the simplifying assumptions $c_{jk}^{a}=0$, thereby suppressing the axial contribution. Suitable values are assigned to $c_{jk}^{v}$, as indicated in each figure. The red line corresponds to the scenario where all three lepton contributions are active within the loop, i.e., $\tau$, $\mu$, and $e$. Conversely, when only a single lepton contribution is activated into the loop, we represent it with purple, green, and blue lines for $e$, $\mu$, and $\tau$, respectively. To contextualize our findings, the grey line denotes the current upper limit, as tabulated in Table 1. Dashed lines are employed to address cases of line overlap, ensuring clarity in the visualization of results.
@@
Our observations demonstrate distinct dominant contributions for each decay process, as depicted in Figures 2(a), 2(b), and 2(c). For the decay $H\to\tau\mu$, the diagrams featuring $\tau$ and $\mu$ within the loop yield the dominant contributions. In the case of $H\to\tau e$, all three diagrams, with $\tau$, $\mu$, and $e$ in the loop, exhibit comparable significance. Conversely, for $H\to\mu e$, the dominant diagrams involve $\mu$ and $e$. This behavior is explained by Eq. (11), which states that $\mathrm{BR}(H\to\ell_{i}\ell_{j})\propto\frac{m_{\ell_{k}}^{2}}{m_{\ell_{i}^{k}}^{2}m_{\ell_{j}^{k}}^{2}}$. The masses in the denominator result from the relationship between the effective couplings $f_{jk}$ and $g_{jk}$ with the coefficients $c_{jk}^{v}$ and $c_{jk}^{a}$ as described in Eq. (10).
@@
By employing the upper limits provided for ${\rm BR}(H\to e\mu\,,e\tau\,,\mu\tau)$ and the flavor-conserving processes ${\rm BR}(H\to ee\,,\mu\mu\,,\tau\tau)$, which are detailed in Table 1, we can discern constraints on the individual coefficients $|c_{ik}^{v(a)}|$. Under the assumption that all $|c_{ik}^{v}|\neq 0~{}{\rm and}~{}|c_{ik}^{a}|=0$, and fixing the parameter $m_{\chi}=m_{\mu}/2$, we derive restrictions for all $|c_{ik}^{v}|$.
@@
It is important to emphasize that these constraint regions remain valid across different values of $m_{\chi}$, as we have shown that ${\rm BR}(H\to\ell_{i}\ell_{j})$ exhibits minimal dependence on $m_{\chi}$. Furthermore, under the assumption that $c_{ik}^{a}\neq 0$ and $c_{ik}^{v}=0$, and keeping $m_{\chi}$ consistent, the resultant constraint regions would remain consistent. This limit on the coefficients $c_{ii}^{v}$ can, in turn, be translated into constraints on the Yukawa couplings, $U(1)_{\chi}$ charges and gauge coupling, and expectation values of the scalar doublets $\phi_{ij}$ in the tree-level model, ensuring the correct couplings with the Higgs are reproduced. Let us finish this section by noting that while previous studies, such as the one in [30], suggest that the most stringent constraints on the coefficients $c_{jk}^{v(a)}$ would likely arise from lepton-flavor violating three-body decays, such as $\mu\to 3e$ and $\tau\to 3\ell$ (with $\ell=e,\mu$), the focus of this work lies elsewhere. Our aim is not to derive the most restrictive bounds on these coefficients but to explore the unique lepton-flavor violating Higgs decays mediated by an ultralight gauge boson. We seek to investigate the phenomenology of these decays in the Higgs sector, providing observables that are sensitive to the mass of the $\chi$ boson, as we will discuss in the next section.